A 100 nC charge is placed at the origin. A -50 nC charge is placed at x= 5 cm and a 200 nC is placed at x= -10 cm.









a. What is the net electrostatic force acting on 100 nC charge?

1
Expert's answer
2022-03-09T10:32:50-0500

Solution.

"q_1=100\sdot10^{-9}C;"

"q_2=-50\sdot10^{-9}C;"

"q_3=200\sdot10^{-9}C;"

"x_2=5cm=0.05m;"

"x_3=-10cm=-0.1m;"

"F-?;"

"F_{21}=k\dfrac{q_2q_1}{r_2^2};"

"F_{21}=9\sdot10^9\dfrac{50\sdot10^{-9}\sdot100\sdot10^{-9}}{0.05^2}=1.8\sdot10^{-2}N;"

"F_{31}=k\dfrac{q_3q_1}{r_3^2};"

"F_{31}=9\sdot10^9\dfrac{200\sdot10^{-9}\sdot100\sdot10^{-9}}{0.1^2}=1.8\sdot10^{-2}N;"

"F=F_{21}-F_{31};"

"F=1.8\sdot10^{-2}-1.8\sdot10^{-2}=0N;"

Answer: "F=0N."